A horizontal beam of unpolarized light is incident on a stack of three polarizing filters with their polarization axes oriented, in sequence, 30◦, 60◦ and 90◦ clockwise from the vertical. The intensity of the light emerging from the stack is measured to be 275 W/m2. What is the intensity of the emerging light (in W/m2) if the middle polarizing filter is removed?
Answer #1

In first case

I = I0 × (1/2) × cos230° × cos230° = 9 I0/ 32 = 275 W/m2

(As angle between succesive filters is 30°

I0 = 275×32/9 = 977.8 W/m2

In second case

I = I0 × (1/2) × cos260° = I0/8 = 977.8/8 = 122.2 W/m2